Brad is nine more than two times as old as Ben. In 15 years the sum of their ages will be 96. How old is Brad now?
Answer by scott8148(6628)

let x=Brad and y=Ben

"Brad is nine more than two times as old as Ben" __ x=2y+9 __ x-9=2y __ (x-9)/2=y

"In 15 years the sum of their ages will be 96" __ (x+15)+(y+15)=96

substituting __ x+15+[(x-9)/2]+15=96 __ 2x+30+x-9+30=192 __ 3x+51=192 __3x=141 __ x=47